OpenStudy (anonymous):

Evaluate the following integral. int_0^a square root a^2-x^2 dx

OpenStudy (anonymous):

that is the area under the circle with radius a, a quarter circle

OpenStudy (anonymous):



OpenStudy (anonymous):

Using geometry you can find the integral (the area ) without doing any calculus, what is the area of a circle with radius `a` ?

OpenStudy (anonymous):

Thank you

OpenStudy (thomas5267):

Or: \[ \begin{align*} &\quad\int_0^a\sqrt{a^2-x^2}\,dx\\ &\qquad x^2=a^2\sin^2(\theta)\\ &\qquad x=a\sin(\theta)\\ &\qquad dx=a\cos(\theta)\,d\theta\\ &=\int_{x=0}^{x=a}\sqrt{a^2-a^2\sin^2(\theta)}\left(a\cos(\theta)\right)\,d\theta\\ &=\int_{x=0}^{x=a}a^2\cos^2(\theta)\,d\theta\\ &=a^2\int_{x=0}^{x=a}\frac{\cos(2\theta)+1}{2}\,d\theta\\ &=a^2\left[\frac{1}{4}\sin(2\theta)+\frac{1}{2}\theta\right]_{x=0}^{x=a}\\ &\qquad x=a\sin(\theta)\\ &\qquad\theta=\arcsin\left(\frac{x}{a}\right)\\ &\qquad x=a\implies \theta=\frac{\pi}{2}\\ &\qquad x=0\implies \theta=0\\ &=a^2\left[\frac{1}{4}\sin(2\theta)+\frac{1}{2}\theta\right]_{0}^{\frac{\pi}{2}}\\ &=a^2\left[\frac{1}{4}\sin(\pi)+\frac{1}{4}\pi-0\right]\\ &=\frac{1}{4}\pi a^2 \text{ as required} \end{align*} \]